# TARDis: Time Attenuated Representation Disentanglement for Incomplete Multi-Modal Tumor Segmentation and Classification

### v1

@@ 3.7 Handling Variable Modality Sequences
To simulate the "missing modality" problem explicitly during training, we employ an aggressive Modality Dropout strategy. Given a subject with a complete set of multi-phase acquisitions, we randomly subset the inputs to a smaller size $N^{\prime}\leq N$ before passing them to the shared encoder. This stochastic dropping prevents the model from relying on any specific combination of phases and forces the shared encoder to extract robust anatomical features from sparse temporal samples.
@@ 4 Results
Figure 4: The distribution of modality combinations in CT datasets. N: Non-Contrast Scan. A: Arterial Phase. V: Portal Venous Phase. D: Delayed Phase. CH: ChangHai Dataset. KiTS19: Kidney Tumor Segmentation 2019 Dataset.
Figure 5: The metrics curves of comparison methods with different modality combinations in CH dataset. N: Non-Contrast Scan. A: Arterial Phase. V: Portal Venous Phase. D: Delayed Phase.
@@
Modality Screening AUC Benign-Malignant AUC 4-Class Diagnosis AUC
N A V D [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis
@@
N A V D [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis
@@
Table 3: Segmentation and screening performance in KiTS19 dataset
@@
[ding2021rfnet] 0.767 0.811 0.835 0.904 0.909 0.933
[zhang2022mmformer] 0.703 0.787 0.774 0.855 0.886 0.902
[liu2023m3ae] 0.790 0.840 0.836 0.940 0.955 0.948
[m2ftrans] 0.321 0.394 0.675 0.735 0.682 0.773
@@
T1 T1CE T2 FLAIR [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis [ding2021rfnet] [zhang2022mmformer] [liu2023m3ae] [m2ftrans] TARDis
@@
Figure 6: The metrics curves of comparison methods with different modality combinations in the BraTS18 dataset. The number on the x-axis refers to the modality combination in Table 4.
@@ 4.1 Datasets
The first dataset is the Climb 4 Kidney Cancer - Kidney and Kidney Tumor Segmentation (C4KC-KiTS) collection [C4KCKITS2019, heller2019kits19, heller2020state], a public subset of the KiTS19 challenge data. We utilized 210 cases from the training set. As shown in Fig. 4, every case includes an arterial phase scan, while a subset of cases also contains non-contrast or delayed phase acquisitions, providing a controlled environment for testing missing-modality performance.
Changhai (CH) Dataset
To validate our method in a large-scale clinical setting, we collected a private multi-phase abdominal CT dataset from Changhai Hospital, Shanghai, China. This extensive cohort contains 2,282 cases comprising 8,065 total scans (Fig. 4). The dataset includes 2,958 annotated kidney lesions spanning 26 distinct histological subtypes. For experimental purposes, these subtypes were categorized into four clinically relevant classes based on surgical management strategies: cysts (2,188), other benign lesions (105), low-grade malignant lesions (450), and high-grade malignant lesions (215). All volumes were spatially registered and preprocessed to facilitate multi-modal tumor diagnosis. Ground truth labels were annotated and iteratively refined by board-certified radiologists.
@@
While TARDis is theoretically grounded in the CT modalities, we extended our evaluation to the BraTS 2018 dataset [menze2014multimodal, bakas2017advancing, bakas2018identifying] to benchmark against existing state-of-the-art (SOTA) incomplete modality frameworks, which are predominantly established on this neuroimaging standard. BraTS18 provides four co-registered MRI modalities per case: T1-weighted (T1), post-contrast T1-weighted (T1CE), T2-weighted (T2), and FLAIR. Annotations include three nested sub-regions: GD-enhancing tumor (ET), peritumoral edema (ED), and the necrotic/non-enhancing tumor core (NCR/NET). Following standard challenge protocols, we evaluated performance on three aggregated regions: Whole Tumor (WT), Tumor Core (TC), and Enhancing Tumor (ET).
@@ 4.4 Comparisons with the State-of-the-Art Methods
We compared TARDis against four state-of-the-art frameworks designed for incomplete multi-modal segmentation: RFNet, mmFormer, M3AE, and M2FTrans. The quantitative results across the CH, C4KC-KiTS, and BraTS18 datasets are summarized in Tables 1, 2, 3, and 4.
On the large-scale CH dataset, TARDis consistently outperformed all baselines across both classification and segmentation tasks. As shown in Table 1, our method achieved an average Screening AUC of 0.979, significantly surpassing the second-best method, M2FTrans (0.937). Notably, in the challenging 4-Class Diagnosis task, TARDis demonstrated remarkable superiority with an AUC of 0.923, whereas competing methods struggled to exceed 0.76. This substantial margin suggests that explicitly disentangling the dynamic tumor enhancement significantly aids in distinguishing subtle histological subtypes.
@@ 4.5 Modality Ablation and Robustness Analysis
In kidney tumor diagnosis, contrast-enhanced phases are clinically critical. Fig. 5 illustrates that baseline methods exhibit jagged performance curves, with deep troughs occurring whenever the ’A’ or ’V’ phases are absent. In contrast, TARDis exhibits a much smoother performance curve. For instance, in the Benign-Malignant classification task, standard methods drop to AUCs below 0.60 when limited to non-contrast data. TARDis, however, leverages the CVAE-based modal-specific path to hallucinate the missing perfusion features, maintaining an AUC above 0.85 even in the absence of contrast media.
The "missing modality" problem is most severe when only a single modality is available. Table 2 highlights that under the single-modality setting (average of N, A, V, D individual performance), TARDis achieves an average Dice of roughly 0.86, whereas RFNet and mmFormer lag significantly behind at approximately 0.74 and 0.63, respectively. This robustness is further corroborated by Fig.6 for the BraTS dataset, where the variance in Dice scores across different combinations is visibly lower for TARDis compared to baselines. This stability implies that our Disentangled Time Attenuated Representation effectively decouples the diagnostic signal from the modality availability, ensuring reliable clinical deployment regardless of the acquisition protocol.
@@ 4.6 Qualitative Analysis and Visualization
To intuitively understand the efficacy of TARDis, we present qualitative segmentation results under the most challenging scenario where only a single modality is available (Fig. 7). In these cases, standard missing-modality frameworks such as RFNet, mmFormer, and M3AE struggle significantly. As observed in the visual comparisons, these baseline methods often produce fragmented segmentations or fail to detect the tumor entirely, particularly when the highly informative contrast-enhanced phases (arterial or venous) are absent. This failure likely stems from their reliance on channel fusion or masking strategies; when a specific channel is missing, the corresponding diagnostic signal is effectively lost, and the model cannot compensate for the missing hemodynamic information. In contrast, TARDis maintains high structural consistency with the ground truth across all single-modality inputs. By leveraging the CVAE-based modal-specific path, our model effectively hallucinates the missing perfusion features conditioned on the available static anatomy and the inferred relative time. This capability ensures that the diagnostic boundaries remain distinct even in non-contrast scans, demonstrating the robustness of our generative approach against extreme data sparsity.
@@ 5 Conclusion
In this paper, we presented TARDis, a unified framework for incomplete multi-modal tumor segmentation and classification that moves beyond standard channel-masking strategies. By grounding our architectural design in the physical principles of CT hemodynamics, we hypothesized that the diagnostic signal is composed of a static anatomical baseline and a dynamic perfusion process. The integration of a dictionary-based anatomical encoder ensures structural consistency across scans, while the time-conditioned CVAE allows for the probabilistic reconstruction of dynamic features, even when specific contrast phases are physically absent. Our comprehensive evaluation across large-scale CT and MRI datasets confirms that this physics-aware approach yields superior generalization compared to existing methods that treat modalities as uncorrelated inputs. Specifically, the qualitative and quantitative improvements observed in single-modality inference scenarios suggest that TARDis can effectively infer missing hemodynamic information from learned latent priors, which validates our hypothesis. Clinically, this work suggests a pathway toward more flexible acquisition protocols. Future work will explore extending this continuous-time modeling to other scenarios like low-dose diagnostic workflows.

### v2

@@ 3.7 Handling Variable Modality Sequences
To simulate the missing modality problem explicitly during training, we employ an aggressive Modality Dropout strategy. Given a subject with a complete set of multi-phase acquisitions, we randomly subset the inputs to a smaller size $N^{\prime}\leq N$ before passing them to the shared encoder. This stochastic dropping prevents the model from relying on any specific combination of phases and forces the shared encoder to extract robust anatomical features from sparse temporal samples.
@@ 4 Results
Figure 4: The distribution of modality combinations in CT datasets. N: Non-Contrast Scan. A: Arterial Phase. V: Portal Venous Phase. D: Delayed Phase. CH: ChangHai Dataset.
Figure 5: The metrics curves of compared methods with different modality combinations in CH dataset. N: Non-Contrast Scan. A: Arterial Phase. V: Portal Venous Phase. D: Delayed Phase.
@@
Modality Screening AUC Benign-Malignant AUC Subtype Diagnosis AUC
N A V D RFNet mmFormer M3AE M2FTrans TARDis RFNet mmFormer M3AE M2FTrans TARDis RFNet mmFormer M3AE M2FTrans TARDis
@@
N A V D RFNet mmFormer M3AE M2FTrans TARDis
@@
Table 3: Segmentation and screening performance in C4KC-KiTS dataset
@@
RFNet 0.767 0.811 0.835 0.904 0.909 0.933
mmFormer 0.703 0.787 0.774 0.855 0.886 0.902
M3AE 0.790 0.840 0.836 0.940 0.955 0.948
M2FTrans 0.321 0.394 0.675 0.735 0.682 0.773
@@
T1 T1CE T2 FLAIR RFNet mmFormer M3AE M2FTrans TARDis RFNet mmFormer M3AE M2FTrans TARDis RFNet mmFormer M3AE M2FTrans TARDis
@@
Figure 6: The metrics curves of compared methods with different modality combinations in the BraTS18 dataset. The number on the x-axis refers to the modality combination in Table 4.
@@ 4.1 Datasets
Changhai (CH) Dataset
To validate our method in a large-scale multi-modal clinical setting, we collected a private multi-phase abdominal CT dataset from Changhai Hospital, Shanghai, China. This extensive cohort contains 2,282 patients comprising 8,065 total scans (Fig. 4). The dataset includes 2,958 annotated kidney lesions spanning 26 distinct histological subtypes. For experimental purposes, these subtypes were categorized into four clinically relevant classes based on surgical management strategies: cysts (2,188), other benign lesions (105), low-grade malignant lesions (450), and high-grade malignant lesions (215). All volumes were spatially registered and preprocessed to facilitate multi-modal tumor diagnosis. Ground truth labels were annotated and iteratively refined by board-certified radiologists.
@@
The first dataset is the Climb 4 Kidney Cancer - Kidney and Kidney Tumor Segmentation (C4KC-KiTS) collection [C4KCKITS2019, heller2019kits19, heller2020state], a public subset of the KiTS19 challenge data. We utilized 210 patients from the training set. As shown in Fig. 4, every patient includes an arterial phase scan, while a subset of patients also contains non-contrast or delayed phase acquisitions, providing a controlled environment for testing missing-modality performance.
@@
While TARDis is theoretically grounded in the CT modalities, we extended our evaluation to the BraTS 2018 dataset [menze2014multimodal, bakas2017advancing, bakas2018identifying] to compare with existing state-of-the-art (SOTA) incomplete modality frameworks, which are predominantly established on this neuroimaging standard. BraTS18 provides four co-registered MRI modalities per patient: T1-weighted (T1), post-contrast T1-weighted (T1CE), T2-weighted (T2), and FLAIR. Annotations include three nested sub-regions: GD-enhancing tumor (ET), peritumoral edema (ED), and the necrotic/non-enhancing tumor core (NCR/NET). Following standard challenge protocols, we evaluated performance on three aggregated regions: Whole Tumor (WT), Tumor Core (TC), and Enhancing Tumor (ET).
@@ 4.4 Comparisons with the State-of-the-Art Methods
We compared TARDis against four state-of-the-art frameworks. The quantitative results across the CH, C4KC-KiTS, and BraTS18 datasets are summarized in Tables 1, 2, 3, and 4. On the large-scale multi-modal CH dataset, TARDis consistently outperformed all baselines across both classification and segmentation tasks. As shown in Table 1, our method achieved an average Screening AUC of 0.979, significantly surpassing the second-best method, M2FTrans (0.937). Notably, in the challenging Subtype Diagnosis task, TARDis demonstrated remarkable superiority with an AUC of 0.923, whereas competing methods struggled to exceed 0.760. This substantial margin suggests that explicitly disentangling the dynamic tumor enhancement significantly aids in distinguishing subtle histological subtypes.
@@ 4.5 Modality Ablation and Robustness Analysis
In kidney tumor diagnosis, contrast-enhanced phases are clinically critical. Fig. 5 illustrates that baseline methods exhibit jagged performance curves, with deep troughs occurring whenever the ’A’ or ’V’ phases are absent. In contrast, TARDis exhibits a much smoother performance curve. For instance, in the Benign-Malignant classification task, standard methods drop to AUCs below 0.60 when limited to non-contrast data. TARDis, however, leverages the CVAE-based modal-specific path to infer the missing perfusion features, maintaining an AUC above 0.85 even in the absence of contrast media.
The missing modality problem is most severe when only a single modality is available. Table 2 highlights that under the single-modality setting (average of N, A, V, D individual performance), TARDis achieves an average Dice of roughly 0.86, whereas RFNet and mmFormer lag significantly behind at approximately 0.74 and 0.63, respectively. This robustness is further corroborated by Fig.6 for the BraTS dataset, where the variance in Dice scores across different combinations is visibly lower for TARDis compared to baselines. This stability implies that our Disentangled Time Attenuated Representation effectively decouples the diagnostic signal from the modality availability, ensuring reliable clinical deployment regardless of the acquisition protocol.
@@ 4.6 Qualitative Analysis and Visualization
To intuitively understand the efficacy of TARDis, we present qualitative segmentation results under the most challenging scenario where only a single modality is available (Fig. 7). In these cases, standard missing-modality frameworks such as RFNet, mmFormer, and M3AE struggle significantly. As observed in the visual comparisons, these baseline methods often produce fragmented segmentations or fail to detect the tumor entirely, particularly when the highly informative contrast-enhanced phases (arterial or venous) are absent. This failure likely stems from their reliance on channel fusion or masking strategies; when a specific channel is missing, the corresponding diagnostic signal is effectively lost, and the model cannot compensate for the missing hemodynamic information. In contrast, TARDis maintains high structural consistency with the ground truth across all single-modality inputs. By leveraging the CVAE-based modal-specific path, our model effectively infers the missing perfusion features conditioned on the available static anatomy and the inferred relative time. This capability ensures that the diagnostic boundaries remain distinct even in non-contrast scans, demonstrating the robustness of our generative approach against extreme data sparsity.
@@ 5 Conclusion
In this paper, we presented TARDis, a unified framework for incomplete multi-modal tumor segmentation and classification that moves beyond standard channel-masking strategies. By grounding our architectural design in the physical principles of CT hemodynamics, we hypothesized that the diagnostic signal is composed of a static anatomical baseline and a dynamic perfusion process. The integration of a dictionary-based anatomical encoder ensures structural consistency across scans, while the time-conditioned HCVAE allows for the probabilistic reconstruction of dynamic features, even when specific contrast phases are physically absent. Our comprehensive evaluation across large-scale multi-modal CT and MRI datasets confirms that this physics-aware approach yields superior generalization compared to existing methods that treat modalities as uncorrelated inputs. Specifically, the qualitative and quantitative improvements observed in single-modality inference scenarios suggest that TARDis can effectively infer missing hemodynamic information from learned latent priors, which validates our hypothesis. Clinically, by relieving the dependency on complete multi-phase series, TARDis paves the way for flexible acquisition protocols. This capability creates a significant opportunity to reduce scan durations and minimize patient radiation exposure, thereby improving overall hospital efficiency and throughput without compromising diagnostic precision. Furthermore, we believe this continuous-time disentanglement paradigm can be generalized beyond segmentation to broader medical challenges, such as low-dose diagnostic and cross-center data harmonization.